# Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers

### v1

@@ Acknowledgements.
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills, and G. Csányi, A foundation model for atomistic materials chemistry (2023), arXiv:2401.00096 [physics.chem-ph] .
- Xie et al. [2021] T. Xie, X. Fu, O.-E. Ganea, R. Barzilay, and T. Jaakkola, Crystal diffusion variational autoencoder for periodic material generation, arXiv preprint arXiv:2110.06197 (2021).
- Wu et al. [2022] L. Wu, C. Gong, X. Liu, M. Ye, and Q. Liu, Diffusion-based molecule generation with informative prior bridges, Advances in Neural Information Processing Systems 35, 36533 (2022).
@@
- Zaidi et al. [2022] S. Zaidi, M. Schaarschmidt, J. Martens, H. Kim, Y. W. Teh, A. Sanchez-Gonzalez, P. Battaglia, R. Pascanu, and J. Godwin, Pre-training via denoising for molecular property prediction, arXiv preprint arXiv:2206.00133 (2022).
- Jia et al. [2024] S. Jia, A. R. Parthasarathy, R. Feng, G. Cong, C. Zhang, and V. Fung, Derivative-based pre-training of graph neural networks for materials property predictions, Digital Discovery (2024).
- Ramakrishnan et al. [2014] R. Ramakrishnan, P. O. Dral, M. Rupp, and O. A. von Lilienfeld, Quantum chemistry structures and properties of 134 kilo molecules, Scientific Data 1, 10.1038/sdata.2014.22 (2014).
- Dral et al. [2015] P. O. Dral, O. A. von Lilienfeld, and W. Thiel, Machine learning of parameters for accurate semiempirical quantum chemical calculations, Journal of chemical theory and computation 11, 2120 (2015).
@@
- Pedregosa et al. [2011] F. Pedregosa, G. Varoquaux, A. Gramfort, V. Michel, B. Thirion, O. Grisel, M. Blondel, P. Prettenhofer, R. Weiss, V. Dubourg, J. Vanderplas, A. Passos, D. Cournapeau, M. Brucher, M. Perrot, and E. Duchesnay, Scikit-learn: Machine learning in Python, Journal of Machine Learning Research 12, 2825 (2011).
@@
- Akiba et al. [2019] T. Akiba, S. Sano, T. Yanase, T. Ohta, and M. Koyama, Optuna: A next-generation hyperparameter optimization framework (2019), arXiv:1907.10902 [cs.LG] .
- Schuld and Killoran [2019] M. Schuld and N. Killoran, Quantum machine learning in feature hilbert spaces, Physical Review Letters 122, 10.1103/physrevlett.122.040504 (2019).
- Kusumoto et al. [2021] T. Kusumoto, K. Mitarai, K. Fujii, M. Kitagawa, and M. Negoro, Experimental quantum kernel trick with nuclear spins in a solid, npj Quantum Information 7, 10.1038/s41534-021-00423-0 (2021).
- Haug et al. [2023] T. Haug, C. N. Self, and M. S. Kim, Quantum machine learning of large datasets using randomized measurements, Machine Learning: Science and Technology 4, 015005 (2023).
- Haug and Kim [2022] T. Haug and M. S. Kim, Natural parametrized quantum circuit, Physical Review A 106, 10.1103/physreva.106.052611 (2022).
- Benedetti et al. [2019] M. Benedetti, E. Lloyd, S. Sack, and M. Fiorentini, Parameterized quantum circuits as machine learning models, Quantum Science and Technology 4, 043001 (2019).
- [76] https://github.com/Qulacs-Osaka/scikit-qulacs.
@@
- Suzuki et al. [2021] Y. Suzuki, Y. Kawase, Y. Masumura, Y. Hiraga, M. Nakadai, J. Chen, K. M. Nakanishi, K. Mitarai, R. Imai, S. Tamiya, T. Yamamoto, T. Yan, T. Kawakubo, Y. O. Nakagawa, Y. Ibe, Y. Zhang, H. Yamashita, H. Yoshimura, A. Hayashi, and K. Fujii, Qulacs: a fast and versatile quantum circuit simulator for research purpose, Quantum 5, 559 (2021).
@@
- Christensen et al. [2017] A. S. Christensen, L. A. Bratholm, F. A. Faber, B. Huang, A. Tkatchenko, K. R. Müller, and O. A. von Lilienfeld, Qml: A python toolkit for quantum machine learning, https://github.com/qmlcode/qml (2017).
- Lopanitsyna et al. [2023] N. Lopanitsyna, G. Fraux, M. A. Springer, S. De, and M. Ceriotti, Modeling high-entropy transition metal alloys with alchemical compression, Physical Review Materials 7, 045802 (2023).
- Willatt et al. [2018] M. J. Willatt, F. Musil, and M. Ceriotti, Feature optimization for atomistic machine learning yields a data-driven construction of the periodic table of the elements, Physical Chemistry Chemical Physics 20, 29661 (2018).
@@
- Li et al. [2022] S. Li, Y. Liu, D. Chen, Y. Jiang, Z. Nie, and F. Pan, Encoding the atomic structure for machine learning in materials science, Wiley Interdisciplinary Reviews: Computational Molecular Science 12, e1558 (2022).
- Behler [2011] J. Behler, Atom-centered symmetry functions for constructing high-dimensional neural network potentials, The Journal of Chemical Physics 134, 074106 (2011).
- Isert et al. [2022] C. Isert, K. Atz, J. Jiménez-Luna, and G. Schneider, Qmugs, quantum mechanical properties of drug-like molecules, Scientific Data 9, 273 (2022).
@@
- Xin et al. [2017] D. Xin, C. A. Sader, U. Fischer, K. Wagner, P.-J. Jones, M. Xing, K. R. Fandrick, and N. C. Gonnella, Systematic investigation of dft-giao 15n nmr chemical shift prediction using b3lyp/cc-pvdz: application to studies of regioisomers, tautomers, protonation states and n-oxides, Organic & Biomolecular Chemistry 15, 928–936 (2017).
@@
- Puzzarini et al. [2009] C. Puzzarini, G. Cazzoli, M. E. Harding, J. Vázquez, and J. Gauss, A new experimental absolute nuclear magnetic shielding scale for oxygen based on the rotational hyperfine structure of h2o17, The Journal of Chemical Physics 131, 10.1063/1.3274062 (2009).
- Wasylishen and Bryce [2002] R. E. Wasylishen and D. L. Bryce, A revised experimental absolute magnetic shielding scale for oxygen, The Journal of Chemical Physics 117, 10061–10066 (2002).
@@
- Rosenau et al. [2018] C. P. Rosenau, B. J. Jelier, A. D. Gossert, and A. Togni, Exposing the origins of irreproducibility in fluorine nmr spectroscopy, Angewandte Chemie International Edition 57, 9528 (2018).
@@
- Frisch et al. [2016] M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman, and D. J. Fox, Gaussian˜16 Revision C.01 (2016), gaussian Inc. Wallingford CT.
- Rupp et al. [2012] M. Rupp, A. Tkatchenko, K.-R. Müller, and O. A. von Lilienfeld, Fast and accurate modeling of molecular atomization energies with machine learning, Physical Review Letters 108, 10.1103/physrevlett.108.058301 (2012).
- Samman et al. [2023] A. M. E. Samman, S. D. Castro, B. Morton, and S. D. Baerdemacker, Transfer learning graph representations of molecules for pka, 13c-nmr, and solubility 10.26434/chemrxiv-2023-hfcm5 (2023).
@@
- Cerezo et al. [2021] M. Cerezo, A. Arrasmith, R. Babbush, S. C. Benjamin, S. Endo, K. Fujii, J. R. McClean, K. Mitarai, X. Yuan, L. Cincio, and P. J. Coles, Variational quantum algorithms, Nature Reviews Physics 3, 625 (2021).
- Mitarai et al. [2018] K. Mitarai, M. Negoro, M. Kitagawa, and K. Fujii, Quantum circuit learning, Physical Review A 98, 10.1103/physreva.98.032309 (2018).
- El-Samman et al. [2023] A. El-Samman, I. Husain, M. Huynh, S. De Castro, B. Morton, G. Acke, and S. De Baerdemacker, Global interpretability and geometry of graph convolu- tional neural networks for chemistry in terms of chemical moieties 10.26434/chemrxiv-2023-zhlmr (2023).
- Elijošius et al. [2024] R. Elijošius, F. Zills, I. Batatia, S. W. Norwood, D. P. Kovács, C. Holm, and G. Csányi, Zero shot molecular generation via similarity kernels, arXiv preprint arXiv:2402.08708 (2024).
Appendix A Distribution of Datasets for Each NMR Chemical Shift Prediction Model
The distributions of the training and test sets sampled from the QM9NMR dataset are shown in Fig. 4. Figure 4.(b) shows that above 5K, the distribution is in good agreement with the overall distribution of the C NMR shielding constants. For the other elemental species, the distributions of the training and test sets were in good agreement with the overall distribution.
Appendix B Validation of Learning Accuracy of NMR Chemical Shift Prediction
Figure 5 illustrates the accuracy of the KRR models trained using the M3GNet GNN-TF descriptor for five elemental species. The MAE values for the NMR shielding constant, for train/test, are as follows: for , 0.0344/0.1767; , 0.1420/2.2798 ppm; , 0.3910/3.4157 ppm; 0.8881/4.9509 ppm; and 0.0864/2.6518 ppm.
The structures of two sets of drug molecules [29] used for validating the constructed model beyond the QM9NMR dataset are depicted in Figs. 6 and 7, respectively. The predicted C NMR shielding constants for each drug molecule using the M3GNet GNN-TF/KRR model are shown in Fig. 8. These predictions are accompanied by the values predicted by the FCHL/KRR model.[29]

### v2

@@ Acknowledgements
- Batatia et al. 2023 I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills and G. Csányi, A foundation model for atomistic materials chemistry, 2023.
- Merchant et al. 2023 A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon and E. D. Cubuk, Nature, 2023, 624, 80–85.
- Batzner et al. 2022 S. Batzner, A. Musaelian, L. Sun, M. Geiger, J. P. Mailoa, M. Kornbluth, N. Molinari, T. E. Smidt and B. Kozinsky, Nature communications, 2022, 13, 2453.
- Riebesell et al. 2024 J. Riebesell, R. E. A. Goodall, P. Benner, Y. Chiang, B. Deng, A. A. Lee, A. Jain and K. A. Persson, Matbench Discovery – A framework to evaluate machine learning crystal stability predictions, 2024.
- Han et al. 2022 J. Han, H. Kang, S. Kang, Y. Kwon, D. Lee and Y.-S. Choi, Physical Chemistry Chemical Physics, 2022, 24, 26870–26878.
- Grimblat et al. 2015 N. Grimblat, M. M. Zanardi and A. M. Sarotti, The Journal of organic chemistry, 2015, 80, 12526–12534.
- Semenov and Krivdin 2020 V. A. Semenov and L. B. Krivdin, Magnetic Resonance in Chemistry, 2020, 58, 56–64.
- Schaefer et al. 1986 T. Schaefer, J. Peeling and G. H. Penner, Canadian journal of chemistry, 1986, 64, 2162–2167.
- Fukaya and Ono 2004 H. Fukaya and T. Ono, Journal of computational chemistry, 2004, 25, 51–60.
- Chen et al. 2013 H. Chen, S. Viel, F. Ziarelli and L. Peng, Chemical Society Reviews, 2013, 42, 7971–7982.
- Yu et al. 2013 J.-X. Yu, R. R. Hallac, S. Chiguru and R. P. Mason, Progress in nuclear magnetic resonance spectroscopy, 2013, 70, 25–49.
- Gerrard et al. 2022 W. Gerrard, C. Yiu and C. P. Butts, Magnetic Resonance in Chemistry, 2022, 60, 1087–1092.
- Krivdin 2023 L. B. Krivdin, Magnetic Resonance in Chemistry, 2023, 61, 507–529.
- Matsuzaki et al. 2024 K. Matsuzaki, S. Hayashi and W. Nakanishi, RSC advances, 2024, 14, 14340–14356.
- Xie et al. 2021 T. Xie, X. Fu, O.-E. Ganea, R. Barzilay and T. Jaakkola, arXiv preprint arXiv:2110.06197, 2021.
- Wu et al. 2022 L. Wu, C. Gong, X. Liu, M. Ye and Q. Liu, Advances in Neural Information Processing Systems, 2022, 35, 36533–36545.
@@
- Zaidi et al. 2022 S. Zaidi, M. Schaarschmidt, J. Martens, H. Kim, Y. W. Teh, A. Sanchez-Gonzalez, P. Battaglia, R. Pascanu and J. Godwin, arXiv preprint arXiv:2206.00133, 2022.
- Jia et al. 2024 S. Jia, A. R. Parthasarathy, R. Feng, G. Cong, C. Zhang and V. Fung, Digital Discovery, 2024.
- Ramakrishnan et al. 2014 R. Ramakrishnan, P. O. Dral, M. Rupp and O. A. Von Lilienfeld, Scientific data, 2014, 1, 1–7.
- Deng et al. 2023 B. Deng, P. Zhong, K. Jun, J. Riebesell, K. Han, C. J. Bartel and G. Ceder, Nature Machine Intelligence, 2023, 5, 1031–1041.
- Eastman et al. 2023 P. Eastman, P. K. Behara, D. L. Dotson, R. Galvelis, J. E. Herr, J. T. Horton, Y. Mao, J. D. Chodera, B. P. Pritchard, Y. Wang et al., Scientific Data, 2023, 10, 11.
- Isert et al. 2022 C. Isert, K. Atz, J. Jiménez-Luna and G. Schneider, Scientific Data, 2022, 9, 273.
- Elijošius et al. 2024 R. Elijošius, F. Zills, I. Batatia, S. W. Norwood, D. P. Kovács, C. Holm and G. Csányi, arXiv preprint arXiv:2402.08708, 2024.
@@
- Pedregosa et al. 2011 F. Pedregosa, G. Varoquaux, A. Gramfort, V. Michel, B. Thirion, O. Grisel, M. Blondel, P. Prettenhofer, R. Weiss, V. Dubourg, J. Vanderplas, A. Passos, D. Cournapeau, M. Brucher, M. Perrot and E. Duchesnay, Journal of Machine Learning Research, 2011, 12, 2825–2830.
@@
- Akiba et al. 2019 T. Akiba, S. Sano, T. Yanase, T. Ohta and M. Koyama, Optuna: A Next-generation Hyperparameter Optimization Framework, 2019.
- Schuld and Killoran 2019 M. Schuld and N. Killoran, Physical review letters, 2019, 122, 040504.
- Kusumoto et al. 2021 T. Kusumoto, K. Mitarai, K. Fujii, M. Kitagawa and M. Negoro, npj Quantum Information, 2021, 7, 94.
- Haug et al. 2023 T. Haug, C. N. Self and M. S. Kim, Machine Learning: Science and Technology, 2023, 4, 015005.
- Haug and Kim 2022 T. Haug and M. Kim, Physical Review A, 2022, 106, 052611.
- Benedetti et al. 2019 M. Benedetti, E. Lloyd, S. Sack and M. Fiorentini, Quantum Science and Technology, 2019, 4, 043001.
- 94 https://github.com/Qulacs-Osaka/scikit-qulacs.
@@
- Suzuki et al. 2021 Y. Suzuki, Y. Kawase, Y. Masumura, Y. Hiraga, M. Nakadai, J. Chen, K. M. Nakanishi, K. Mitarai, R. Imai, S. Tamiya, T. Yamamoto, T. Yan, T. Kawakubo, Y. O. Nakagawa, Y. Ibe, Y. Zhang, H. Yamashita, H. Yoshimura, A. Hayashi and K. Fujii, Quantum, 2021, 5, 559.
@@
- Christensen et al. 2017 A. S. Christensen, L. A. Bratholm, F. A. Faber, B. Huang, A. Tkatchenko, K. R. Müller and O. A. von Lilienfeld, QML: A Python Toolkit for Quantum Machine Learning, https://github.com/qmlcode/qml, 2017.
- Lopanitsyna et al. 2023 N. Lopanitsyna, G. Fraux, M. A. Springer, S. De and M. Ceriotti, Physical Review Materials, 2023, 7, 045802.
- Willatt et al. 2018 M. J. Willatt, F. Musil and M. Ceriotti, Physical Chemistry Chemical Physics, 2018, 20, 29661–29668.
@@
- Li et al. 2022 S. Li, Y. Liu, D. Chen, Y. Jiang, Z. Nie and F. Pan, Wiley Interdisciplinary Reviews: Computational Molecular Science, 2022, 12, e1558.
- Behler 2011 J. Behler, The Journal of Chemical Physics, 2011, 134, 074106.
@@
- Xin et al. 2017 D. Xin, C. A. Sader, U. Fischer, K. Wagner, P.-J. Jones, M. Xing, K. R. Fandrick and N. C. Gonnella, Organic & Biomolecular Chemistry, 2017, 15, 928–936.
@@
- Puzzarini et al. 2009 C. Puzzarini, G. Cazzoli, M. E. Harding, J. Vázquez and J. Gauss, The Journal of Chemical Physics, 2009, 131, 234304.
- Wasylishen and Bryce 2002 R. E. Wasylishen and D. L. Bryce, The Journal of Chemical Physics, 2002, 117, 10061–10066.
@@
- Rosenau et al. 2018 C. P. Rosenau, B. J. Jelier, A. D. Gossert and A. Togni, Angewandte Chemie International Edition, 2018, 57, 9528–9533.
- Adamo and Barone 1998 C. Adamo and V. Barone, The Journal of chemical physics, 1998, 108, 664–675.
- Ditchfield 1972 R. Ditchfield, The Journal of Chemical Physics, 1972, 56, 5688–5691.
- Stephens et al. 1994 P. J. Stephens, F. J. Devlin, C. F. Chabalowski and M. J. Frisch, The Journal of physical chemistry, 1994, 98, 11623–11627.
@@
- Frisch et al. 2016 M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman and D. J. Fox, Gaussian˜16 Revision C.01, 2016, Gaussian Inc. Wallingford CT.
- Rupp et al. 2012 M. Rupp, A. Tkatchenko, K.-R. Müller and O. A. Von Lilienfeld, Physical review letters, 2012, 108, 058301.
- El-Samman et al. 2023 A. M. El-Samman, S. De Castro, B. Morton and S. De Baerdemacker, Canadian Journal of Chemistry, 2023.
@@
- Cerezo et al. 2021 M. Cerezo, A. Arrasmith, R. Babbush, S. C. Benjamin, S. Endo, K. Fujii, J. R. McClean, K. Mitarai, X. Yuan, L. Cincio and P. J. Coles, Nature Reviews Physics, 2021, 3, 625–644.
- Mitarai et al. 2018 K. Mitarai, M. Negoro, M. Kitagawa and K. Fujii, Physical Review A, 2018, 98, 032309.
- El-Samman et al. 2024 A. M. El-Samman, I. A. Husain, M. Huynh, S. De Castro, B. Morton and S. De Baerdemacker, Digital Discovery, 2024, 3, 544–557.